# Positive maps and extendibility hierarchies from copositive matrices

### v1

@@ 4. Lifting copositivity to pairwise copositivity
In this section, we study the membership problem for the cone $\operatorname{\mathsf{PCOP}}$ in the case of several classes of pairs of matrices. Many of the results in this section are focused on establishing an equivalence with the membership problem for the cone of copositive matrices $\operatorname{\mathsf{COP}}$. Another motivation is the generalization of Example 3.8 to pairs where the first matrix has constant elements off the diagonal, while the second matrix is arbitrarily chosen.
More precisely, recall the following observation from Item 2 from Proposition 3.5:
$(A,B)\in\operatorname{\mathsf{PCOP}}_{n}\implies A+A^{\top}+2\operatorname{Re}(\mathring{B})\in\operatorname{\mathsf{COP}}_{n}.$ (12)
We state next the main result of this section, a converse to Eq. 12 for quite general pairs of matrices.
@@
Suppose $A,N\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{C})$ satisfies $\operatorname{diag}(N)=\operatorname{diag}(A)$ and $N,N-A\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$. Then
$A\in\operatorname{\mathsf{COP}}_{n}\iff(N,A-\mathring{N})\in\operatorname{\mathsf{PCOP}}_{n}.$
In particular, denoting $(A_{+})_{ij}:=\max(A_{ij},0)$ and $(A_{-})_{ij}:=-\min(A_{ij},0)$, we have
$A\in\operatorname{\mathsf{COP}}_{n}\iff(A_{+},\operatorname{diag}(A)-A_{-})\in\operatorname{\mathsf{PCOP}}_{n}.$
In a similar vein, for $)\lambda\geq\max(0,\max(\mathring${A})), we have
$A\in\operatorname{\mathsf{COP}}_{n}\iff(\operatorname{diag}(A)+\lambda\mathring{J}_{n},A-\lambda\mathring{J}_{n})\in\operatorname{\mathsf{PCOP}}_{n}.$
@@
[…]
for all $v,w\in\mathbb{C}^{n}$. Note that that the left hand side is invariant under the entrywise rotation $v,w\mapsto v\odot u,w\odot u^{\prime}$ for $u,u^{\prime}\in\mathbb{T}^{n}$, while the right hand side is always less than $\langle|v|\odot|w|\big|N-A\big||v|\odot|w|\rangle$ from the condition $N-A\in\operatorname{\mathsf{EWP}}_{n}$. Consequently, it suffices to show Eq. 13 for all $v,w\in\mathbb{R}_{+}^{n}$. However, this is now straightforward from the condition $\langle v\odot w|A|v\odot w\rangle\geq 0$ (since $A\in\operatorname{\mathsf{COP}}_{n}$) and
$\langle v\odot v|N|w\odot w\rangle-\langle v\odot w|N|v\odot w\rangle=\sum_{1\leq i<j\leq n}N_{ij}(v_{i}w_{i}-v_{j}w_{j})^{2}\geq 0.$
Now the second and third assertions simply follows from the first statement with $N=A_{+}$ (note that $\operatorname{diag}(A_{+})=\operatorname{diag}(A)$, $\operatorname{diag}(A_{-})=0$, and $A=A_{+}-A_{-}$) and $N=\operatorname{diag}(A)+\lambda\mathring{J}$, respectively. ∎
@@
An alternative way to state the theorem above is: if $A,B\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$ are such that $\operatorname{diag}(A)=\operatorname{diag}(B)$, $A\in\operatorname{\mathsf{EWP}}_{n}$, $\mathring{B}\in\operatorname{\mathsf{EWP}}_{n}$, and $A+\mathring{B}\in\operatorname{\mathsf{COP}}_{n}$, then $(A,B)\in\operatorname{\mathsf{PCOP}}_{n}$.
Remark 4.3.
Recall the Horn matrix $H$ defined in Eq. 3. As we shall later see in Proposition 5.4, $(\operatorname{diag}(H)+\lambda\mathring{J}_{5},H-\lambda\mathring{J}_{5})\in\operatorname{\mathsf{PCOP}}_{n}$ (if and) only if $\lambda\geq 1$. This implies that the lower bound $)\lambda\geq\max(0,\max(\mathring${A})) (see Item 5 in Proposition 3.5) in the previous theorem is optimal as a multiple of $)\max(0,\max(\mathring${A})).
[…]
@@ 5. Pairwise decomposability
We characterize the decomposability property of (conjugate) diagonal unitary covariant maps.
@@
((1) $\implies$ (2)) Assume $\Phi^{\mathsf{DUC}}_{{A},{B}}$ is a decomposable linear map, $\Phi^{\mathsf{DUC}}_{{A},{B}}=\mathcal{E}_{1}+\mathcal{E}_{2}$ where $\mathcal{E}_{1},\mathcal{E}_{2}$ are (resp.) completely positive and completely copositive maps. We use the DUC-twirling map
@@
on both sides, and we have
@@
for CP map $\Phi^{\mathsf{DUC}}_{{A_{1}},{B_{1}}}$ and a co-CP map $\Phi^{\mathsf{DUC}}_{{A_{2}},{B_{2}}}$. Now it is straightforward to verify that $B^{(i)}=B_{i}$, $i=1,2$, satisfy the conditions above. For the reverse direction ((2) $\Longrightarrow$ (1)), we construct a decomposition
@@
Below, the implications in the first row can be either verified by hand from the definitions, or obtained from Theorem 3.4, Proposition 2.14, and Proposition 5.2 from the obvious implications in the second row.
$(A,B)\in\operatorname{\mathsf{PCP}}_{n}$ $\implies$ $(A,B)\in\operatorname{\mathsf{CLDUI}}^{+}_{n} \displaystyle\implies$ $(A,B)\in\operatorname{\mathsf{PDEC}}_{n}$ $\implies$ $(A,B)\in\operatorname{\mathsf{PCOP}}_{n}$
@@
$\Phi^{\mathsf{(C)DUC}}_{{A},{B}}\text{ is EB}$ $\implies$ $\Phi^{\mathsf{(C)DUC}}_{{A},{B}}\text{ is (co-)CP}$ $\implies$ $\Phi^{\mathsf{(C)DUC}}_{{A},{B}}\text{ is dec.}$ $\implies$ $\Phi^{\mathsf{(C)DUC}}_{{A},{B}}\text{ is positive.}$
We start with an example generalizing Example 3.8, obtained by taking $A=\mathring{N}$, $B^{(1)}=0$, and $B^{(2)}=\pm\mathring{N}$ in Definition 5.1.
@@
We have the following strengthening of Item 2 of Proposition 3.5.
@@
If $(A,B)\in\operatorname{\mathsf{PDEC}}_{n}$, then $(A+\mathring{B})+(A+\mathring{B})^{\top}=A+A^{\top}+2\operatorname{Re}(\mathring{B})\in\mathsf{SPN}_{n}$.
@@
We can write $B=B^{(1)}+B^{(2)}$ satisfying the conditions listed in Definition 5.1. Then, we have,
@@
where the last inequality is due to AM-GM. Consequently, we can show the following inclusion.
@@
Let $A\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$ and let $N\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$ with $\operatorname{diag}(A)=\operatorname{diag}(N)$, and suppose $N_{ij}\geq\frac{1}{2}A_{ij}+\frac{1}{4}(A_{ii}+A_{jj})$ for all $i\neq j\in[n]$. Then, we have
$A\in\mathsf{SPN}_{n}\quad\iff\quad(N,A-\mathring{N})\in\operatorname{\mathsf{PDEC}}_{n}.$
Furthermore, if $A$ is positive semidefinite (resp. entrywise non-negative), then the condition $N\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$ (resp. $N\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$ such that $N-A/2\in\operatorname{\mathsf{EWP}}_{n}$) is enough to obtain $(N,A-\mathring{N})\in\operatorname{\mathsf{PDEC}}_{n}$.
@@
The condition $(N,A-\mathring{N})\in\operatorname{\mathsf{PDEC}}_{n}$ implies $A\in\mathsf{SPN}_{n}$ from Proposition 5.5. For the converse, let us assume $A\in\mathsf{SPN}_{n}$ and write $A=A^{(1)}+A^{(2)}$ where $A^{(1)}\in\mathsf{PSD}^{\mathbb{R}}_{n}$ and $A^{(2)}\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$. Then
@@
$(N,A-\mathring{N})=(\operatorname{diag}(A^{(1)}),A^{(1)})+(\operatorname{diag}(A^{(2)})+\mathring{N},A^{(2)}-\mathring{N}),\in\operatorname{\mathsf{PDEC}}_{n}$
@@
$A_{ij}^{(2)}\leq A_{ij}+|A_{ij}^{(1)}|\leq A_{ij}+\sqrt{A_{ii}^{(1)}A_{jj}^{(1)}}\leq A_{ij}+\frac{A_{ii}^{(1)}+A_{jj}^{(1)}}{2}\leq A_{ij}+\frac{A_{ii}+A_{jj}}{2}.$
Therefore, $N-A^{(2)}/2\in\operatorname{\mathsf{EWP}}_{n}$ is satisfied whenever $N_{ij}\geq\frac{1}{2}A_{ij}+\frac{1}{4}(A_{ii}+A_{jj})$ for all $i\neq j$. ∎
@@
The proof above actually shows that, if $A=A^{(1)}+A^{(2)}\in\mathsf{SPN}_{n}$ with $A^{(1)}\in\mathsf{PSD}^{\mathbb{R}}_{n}$ and $A^{(2)}\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$, then
@@
(i.e. $W$ is decomposable), where $\Pi_{S}:=\frac{1}{2}(I_{d^{2}}+F_{d})$ is the projection onto the symmetric subspace $\mathbb{C}^{n}\vee\mathbb{C}^{n}\subset(\mathbb{C}^{n})^{\otimes 2}$. Indeed, the relations from [SN21, Proposition 4.3] imply that
$\Pi_{S}X^{\mathsf{LDUI}}_{{M},{N}}\Pi_{S}=X^{\mathsf{LDUI}}_{{(M+N)/2},{(M+N)/2}}$
whenever $M,N\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$ with $\operatorname{diag}(M)=\operatorname{diag}(N)$.
The following corollary allows us to provide a $(n^{2}-n)/2$-dimensional region of positive non-decomposable linear maps within the classes of DUC and CDUC linear maps.
@@
If $A\in\operatorname{\mathsf{COP}}_{n}\setminus\mathsf{SPN}_{n}$, then both the maps $\Phi^{\mathsf{(C)DUC}}_{{N},{A-\mathring{N}}}$ are positive and non-decomposable for all $N\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$ such that $\operatorname{diag}(N)=\operatorname{diag}(A)$ and $N_{ij}\geq\max(A)$ for all $i>j$. Moreover, if $P\in\mathsf{DNN}_{n}$ with $\langle P,A\rangle<0$, then the Dicke state $X^{\mathsf{LDUI}}_{{P},{P}}$ is PPT entangled with
$\Phi^{\mathsf{DUC}}_{{N},{A-\mathring{N}}}(X^{\mathsf{LDUI}}_{{P},{P}})\not\geq 0.$

### v4

@@ 4. Lifting copositivity to pairwise copositivity
In this section, we study the membership problem for the cone $\operatorname{\mathsf{PCOP}}$ in the case of several classes of pairs of matrices. Many of the results in this section are focused on establishing an equivalence with the membership problem for the cone of copositive matrices $\operatorname{\mathsf{COP}}$. Another motivation is the generalization of Example 3.8 to pairs where we perturb the second component by small amount of matrix, which turns out to be copositive.
More precisely, we apply Corollary 3.10 to obtain the following main result of this section, an efficient method to generate a wide family of pairs in $\operatorname{\mathsf{PCOP}}_{n}$ lifted from a single copositive matrix.
@@
Suppose $M,N\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$ satisfy $\operatorname{diag}(N)=\operatorname{diag}(M)$, $N\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$, and $N_{ij}\geq\frac{1}{2}M_{ij}$ for all $i\neq j\in[n]$. Then
$M\in\operatorname{\mathsf{COP}}_{n}\iff(N,M-\mathring{N})\in\operatorname{\mathsf{PCOP}}_{n}.$
In particular, denoting $(M_{+})_{ij}:=\max(M_{ij},0)$ and $(M_{-})_{ij}:=-\min(M_{ij},0)$, we have
$M\in\operatorname{\mathsf{COP}}_{n}$ $\iff\frac{1}{2}(M_{+}+\operatorname{diag}(M),M_{+}-2M_{-}+\operatorname{diag}(M))\in\operatorname{\mathsf{PCOP}}_{n}$
$\iff(M_{+},\operatorname{diag}(M)-M_{-})\in\operatorname{\mathsf{PCOP}}_{n}.$
In a similar vein, for all $\lambda\geq\frac{1}{2}\max(0,\max(\mathring{M}))$, we have
$M\in\operatorname{\mathsf{COP}}_{n}\iff(\operatorname{diag}(M)+\lambda\mathring{J}_{n},M-\lambda\mathring{J}_{n})\in\operatorname{\mathsf{PCOP}}_{n}.$
@@
$N+N^{\top}+2\operatorname{Re}(\mathring{M}-\mathring{N})=2M\in\mathsf{COP}_{n}.$
For the direction ($\Longrightarrow$), let $O\in D\mathcal{O}_{n}$ and let us decompose $O=O_{+}-O_{-}$ such that both $O_{\pm}$ are diagonal matrices with entries either 0 or $1$ and $O_{+}O_{-}=0$. Also, note that $O_{+}+O_{-}=I_{n}$. Then since $2N-M\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$, we have the following decomposition into sum of copositive matrices
$N+O(\mathring{M}-\mathring{N})O$ $=(N-ONO)+OMO=2(O_{+}NO_{-}+O_{-}NO_{+})+OMO$
$=\underbrace{\big(O_{+}(2N-M)O_{-}+O_{-}(2N-M)O_{+}\big)}_{\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}\subseteq\operatorname{\mathsf{COP}}_{n}}+\underbrace{O_{+}MO_{+}}_{\operatorname{\mathsf{COP}}_{n}}+\underbrace{O_{-}MO_{-}}_{\operatorname{\mathsf{COP}}_{n}}\in\operatorname{\mathsf{COP}}_{n}.$
Since this is true for all $O\in D\mathcal{O}_{n}$, we have by Corollary 3.10 that $(N,M-\mathring{N})\in\operatorname{\mathsf{PCOP}}_{n}$.
Now the second and third assertions simply follows from the first statement with $N=\frac{1}{2}(M_{+}+\operatorname{diag}(M))$, $N=M_{+}$ (note that $\operatorname{diag}(M_{+})=\operatorname{diag}(M)$, $\operatorname{diag}(M_{-})=0$, and $M=M_{+}-M_{-}$), and $N=\operatorname{diag}(M)+\lambda\mathring{J}$, respectively. ∎
@@
The first statement of Theorem 4.1 can be interpreted as follows: for any fixed copositive matrix $M$, one can generate the family
$\mathcal{F}_{M}=\big\{(N,M-\mathring{N}):N\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}},\;\operatorname{diag}(N)=\operatorname{diag}(M),\text{ and }N_{ij}\geq\frac{1}{2}M_{ij}\;\forall\,i\neq j\big\}$
of pairs in $\operatorname{\mathsf{PCOP}}_{n}$. In other words, any pair $(N,M-\mathring{N})\in\mathcal{F}_{M}$ defines positive linear maps $\Phi^{\mathsf{(C)DUC}}_{{N},{M-\mathring{N}}}$. Note that this family is both simple and rich, in the sense that it allows for $(n^{2}-n)/2$ degrees of freedom in the choice of the upper off-diagonal entries $(N_{ij})_{1\leq i<j\leq n}$. In the remainder of the paper, we will show that this construction indeed generates a broad class of positive indecomposable linear maps and entanglement witnesses capable of detecting entanglement in PPT states that are arbitrarily close to separable ones.
Example 4.3.
For the Horn matrix $H$ defined in Eq. 3, the last assertion in Theorem 4.1 implies that
$(I_{5}+\lambda\mathring{J}_{5},H-\lambda\mathring{J}_{5})\in\operatorname{\mathsf{PCOP}}_{5}$ whenever $\lambda\geq\frac{1}{2}$.
We show that the condition $\lambda\geq\frac{1}{2}$ is also necessary (see also Example 5.8), which implies that the lower bound $\lambda\geq\frac{1}{2}\max(0,\max(\mathring{M}))$ in the previous theorem is tight as a multiple of $\max(0,\max(\mathring{M}))$. Indeed, for $t\geq 0$ we may choose $v=(1,0,t,t,0)^{\top}$ and $w=(-1,0,t,t,0)^{\top}$ to compute that
$\langle v^{\odot 2}\big|I_{5}+\lambda\mathring{J}_{5}\big|w^{\odot 2}\rangle+\langle v\odot w|\mathring{H}-\lambda\mathring{J}_{5}|v\odot w\rangle=1+4t^{2}(2\lambda-1).$
Thus, the condition $(I_{5}+\lambda\mathring{J}_{5},H-\lambda\mathring{J}_{5})\in\operatorname{\mathsf{PCOP}}_{5}$ implies that $2\lambda-1\geq-\frac{1}{4t^{2}}$ for all $t>0$. By taking $t\to\infty$, we obtain $\lambda\geq 1/2$.
On the other hand, the bound $\lambda\geq\frac{1}{2}\max(0,\max(\mathring{M}))$ from Theorem 4.1 is not tight in another perspective: if $M\in\mathsf{PSD}^{\mathbb{R}}_{n}$, then $(\operatorname{diag}(M)+\lambda\mathring{J},M-\lambda\mathring{J})\in\operatorname{\mathsf{PCOP}}_{n}$ for all $\lambda\geq 0$ (see Item 6 in Proposition 3.5) while $\max{(\mathring{M})}>0$ whenever $M$ has positive off-diagonal elements.
@@ 5. Pairwise decomposability
Note from the definition that
$(A,B)\in\operatorname{\mathsf{PDEC}}_{n}\iff(\widetilde{A},B)=\big((A\odot A^{\top})^{\odot 1/2},B\big)\in\operatorname{\mathsf{PDEC}}_{n},$ (16)
which is not true in general for $(A,B)\in\operatorname{\mathsf{PCOP}}_{n}$ (we refer to Section 6).
We first characterize the decomposability property of (conjugate) diagonal unitary covariant maps.
@@
[…]
@@
on both sides, and we have by [PJPY24, Proposition 2.1],
@@
[…]
@@
Below, the implications in the first row can be either verified by hand from the definitions, or obtained from Proposition 3.4, 2.14, and 5.2 from the obvious implications in the second row.
$(A,B)\in\operatorname{\mathsf{PCP}}_{n}$ $\implies$ $\begin{cases}(A,B)\in\operatorname{\mathsf{LDUI}}^{+}_{n}\text{ or}\\ (A,B)\in\operatorname{\mathsf{CLDUI}}^{+}_{n}\end{cases} \displaystyle\implies$ $(A,B)\in\operatorname{\mathsf{PDEC}}_{n}$ $\implies$ $(A,B)\in\operatorname{\mathsf{PCOP}}_{n}$
@@
$\Phi^{\mathsf{(C)DUC}}_{{A},{B}}\text{ is EB}$ $\implies$ $\begin{cases}\Phi^{\mathsf{CDUC}}_{{A},{B}}\text{ is CP or}\\ \Phi^{\mathsf{DUC}}_{{A},{B}}\text{ is CP}\end{cases}$ $\implies$ $\Phi^{\mathsf{(C)DUC}}_{{A},{B}}\text{ is dec.}$ $\implies$ $\Phi^{\mathsf{(C)DUC}}_{{A},{B}}\text{ is positive.}$
We start with an example directly strengthening Example 3.8, obtained by taking $A=\mathring{N}$, $B^{(1)}=0$, and $B^{(2)}=\pm\mathring{N}$ in Definition 5.1.
@@
We have the following analogue of Item 3 of Proposition 3.5.
@@
If $(A,B)\in\operatorname{\mathsf{PDEC}}_{n}$, then $\widetilde{A}+{\rm Re}(\mathring{B})\in\mathsf{SPN}_{n}$.
@@
Note that the condition $\widetilde{A}+{\rm Re}(\mathring{B})\in\mathsf{SPN}_{n}$ above also implies the direct analogue of Proposition 3.5, Item 3:
[…]
since $(A+A^{\top})/2-\widetilde{A}\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$.
@@
Let $M\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$ and let $N\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$ with $\operatorname{diag}(M)=\operatorname{diag}(N)$, and suppose $N_{ij}\geq\frac{1}{2}M_{ij}+\frac{1}{4}(M_{ii}+M_{jj})$ for all $i\neq j\in[n]$. Then, we have
$M\in\mathsf{SPN}_{n}\quad\iff\quad(N,M-\mathring{N})\in\operatorname{\mathsf{PDEC}}_{n}.$
Furthermore, if $M$ is positive semidefinite (resp. entrywise non-negative), then the condition $N\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$ (resp. $N\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$ such that $N-M/2\in\operatorname{\mathsf{EWP}}_{n}$) is enough to obtain $(N,M-\mathring{N})\in\operatorname{\mathsf{PDEC}}_{n}$.
@@
The condition $(N,M-\mathring{N})\in\operatorname{\mathsf{PDEC}}_{n}$ implies $M\in\mathsf{SPN}_{n}$ from Proposition 5.5. For the converse, let us assume $M\in\mathsf{SPN}_{n}$ and write $M=A^{(1)}+A^{(2)}$ where $A^{(1)}\in\mathsf{PSD}^{\mathbb{R}}_{n}$ and $A^{(2)}\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$. Then
@@
$(N,M-\mathring{N})=(\operatorname{diag}(A^{(1)}),A^{(1)})+(\operatorname{diag}(A^{(2)})+\mathring{N},A^{(2)}-\mathring{N}),\in\operatorname{\mathsf{PDEC}}_{n}$
@@
$A_{ij}^{(2)}\leq M_{ij}+|A_{ij}^{(1)}|\leq M_{ij}+\sqrt{A_{ii}^{(1)}A_{jj}^{(1)}}\leq M_{ij}+\frac{A_{ii}^{(1)}+A_{jj}^{(1)}}{2}\leq M_{ij}+\frac{M_{ii}+M_{jj}}{2}.$
Therefore, $N-A^{(2)}/2\in\operatorname{\mathsf{EWP}}_{n}$ is satisfied whenever $N_{ij}\geq\frac{1}{2}M_{ij}+\frac{1}{4}(M_{ii}+M_{jj})$ for all $i\neq j$. ∎
@@
The proof above actually shows that, if $M=A^{(1)}+A^{(2)}\in\mathsf{SPN}_{n}$ with $A^{(1)}\in\mathsf{PSD}^{\mathbb{R}}_{n}$ and $A^{(2)}\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$, then
@@
(i.e. $W$ is decomposable), where $\Pi_{S}:=\frac{1}{2}(I_{d^{2}}+F_{d})=\frac{1}{2}(I_{d^{2}}+\sum_{i,j=1}^{d}|ij\rangle\langle ji|)$ is the projection onto the symmetric subspace $\mathbb{C}^{n}\vee\mathbb{C}^{n}\subset(\mathbb{C}^{n})^{\otimes 2}$. Indeed, the relations from [SN21, Proposition 4.3] imply that
$\Pi_{S}X^{\mathsf{LDUI}}_{{A},{B}}\Pi_{S}=X^{\mathsf{LDUI}}_{{(A+B)/2},{(A+B)/2}}$
whenever $A,B\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$ with $\operatorname{diag}(A)=\operatorname{diag}(B)$.
Example 5.8.
For the matrix
$\widetilde{H}:=\begin{bmatrix}1&-1&1&1\\ -1&1&-1&1\\ 1&-1&1&-1\\ 1&1&-1&1\\ \end{bmatrix}=|v\rangle\langle v|+\begin{bmatrix}0&0&0&2\\ 0&0&0&0\\ 0&0&0&0\\ 2&0&0&0\\ \end{bmatrix}\in\mathsf{SPN}_{4},\quad v=(1,-1,1,-1)^{\top},$
Theorem 5.6 implies that $(I_{4}+\lambda\mathring{J}_{4},\widetilde{H}-\lambda\mathring{J}_{4})\in\operatorname{\mathsf{PDEC}}_{4}$ if $\lambda\geq 1$. We claim that this condition is also necessary. Note that $\widetilde{H}$ is a principal $4\times 4$ submatrix of the Horn matrix $H$, and thus the argument in Example 4.3 implies
$(I_{5}+\lambda\mathring{J}_{5},{H}-\lambda\mathring{J}_{5})\in\operatorname{\mathsf{PCOP}}_{5}\;\;\forall\,\lambda\geq\frac{1}{2}\implies(I_{4}+\lambda\mathring{J}_{4},\widetilde{H}-\lambda\mathring{J}_{4})\in\operatorname{\mathsf{PCOP}}_{4}\;\;\forall\,\lambda\geq\frac{1}{2}.$
In particular, we have $(I_{4}+\lambda\mathring{J}_{4},\widetilde{H}-\lambda\mathring{J}_{4})\in\operatorname{\mathsf{PCOP}}_{4}\setminus\operatorname{\mathsf{PDEC}}_{4}$ in the interval $\lambda\in[1/2,1)$, which leads to the positive indecomposable linear maps $\Phi^{\mathsf{(C)DUC}}_{{I_{4}+\lambda\mathring{J}_{4}},{\widetilde{H}-\lambda\mathring{J}_{4}}}$.
To show the claim, suppose $\widetilde{H}-\lambda\mathring{J}_{4}=B^{(1)}+B^{(2)}$ where $B^{(1,2)}$ satisfy the conditions in Definition 5.1, i.e., $B^{(2)}\in\mathcal{M}^{\mathrm{sa}}_{4}(\mathbb{C})$, $B_{ii}^{(2)}\geq 0$, $|B^{(2)}_{ij}|\leq\lambda$, for $i\neq j$, and
$B^{(1)}=\begin{bmatrix}1&-1-\lambda&1-\lambda&1-\lambda\\ -1-\lambda&1&-1-\lambda&1-\lambda\\ 1-\lambda&-1-\lambda&1&-1-\lambda\\ 1-\lambda&1-\lambda&-1-\lambda&1\\ \end{bmatrix}-B^{(2)}\in\mathsf{PSD}^{\mathbb{C}}_{4}.$
First, the psd condition for each of $2\times 2$ principal submatrices of $B^{(1)}$ implies that $\operatorname{diag}(B^{(2)})=0$ and $B^{(2)}_{ij}=-\lambda$ for all $i,j\in[4]$ with |ij|=1|i-j|=1. Next, the condition for $3\times 3$ principal submatrix
$\begin{bmatrix}1&-1&1-\lambda-B^{(2)}_{13}\\ -1&1&-1\\ 1-\lambda-\overline{B^{(2)}_{13}}&-1&1\end{bmatrix}\geq 0\implies\begin{vmatrix}1&-1&1-\lambda-B^{(2)}_{13}\\ -1&1&-1\\ 1-\lambda-\overline{B^{(2)}_{13}}&-1&1\end{vmatrix}=-|\lambda+B^{(2)}_{13}|^{2}\geq 0$
forces that $B^{(2)}_{13}=B^{(2)}_{31}=-\lambda$, and similarly $B^{(2)}_{24}=B^{(2)}_{42}=-\lambda$. Finally, the last condition $B^{(1)}\geq 0$ would force that $1-\lambda-B^{(2)}_{14}=-1$, i.e., $B^{(2)}_{14}=2-\lambda$. Since $|B^{(2)}_{14}|\leq\lambda$, this implies $\lambda\geq 1$.
Finally, the following corollary allows us to provide a $(n^{2}-n)/2$-dimensional region of positive indecomposable linear maps within the classes of DUC and CDUC linear maps.
@@
If $M\in\operatorname{\mathsf{COP}}_{n}\setminus\mathsf{SPN}_{n}$, then both the maps $\Phi^{\mathsf{(C)DUC}}_{{N},{M-\mathring{N}}}$ are positive and non-decomposable for all $N\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$ such that $\operatorname{diag}(N)=\operatorname{diag}(M)$ and $N_{ij}\geq\frac{1}{2}M_{ij}$ for all $i>j$. Moreover, if $P\in\mathsf{DNN}_{n}$ with $\langle P,M\rangle<0$, then the Dicke state $X^{\mathsf{LDUI}}_{{P},{P}}$ is PPT entangled with
$\Phi^{\mathsf{CDUC}}_{{N},{M-\mathring{N}}}(X^{\mathsf{LDUI}}_{{P},{P}})\not\geq 0.$